OpenStudy (anonymous):

How do I solve Q4 of H8P3 - Memory "Time to Charge Capacitance Q2 to VOH"? I tried using the first order dynamics equation: VOH = VG2max + (VOL-VG2max) * exp(-t/TAU) with TAU = R*CGS and R = RON + RPU || ROFF but I get t = 122.669 ps which seems to be wrong.

OpenStudy (anonymous):

I did the same and get the same result....

OpenStudy (anonymous):

at least no error in my numbers but rather in my formula. I'll take a depper look again by tomorrow.

OpenStudy (anonymous):

if i find de solution e post

OpenStudy (anonymous):

VS=5.0V, VOH=3.5V, VIH=3.0V, VIL=0.9V, VOL=0.5V The gate-source capacitance CGS=3.5fF and VT=1.0V. RON=1900.0Ω, ROFF=105.0MΩ and RPU=12.0kΩ. can you tell me why this is wrong? (3.5-4)/(0.5-4) = exp(-t/τ) exp(-t/τ) = 0.142857142857 -t/τ = ln(0.142857142857) = -1.94591014906 t = T * 1.94591014906 R = 1900 + ((12000 * 105000000)/(12000 + 105000000)) = 13898.6287281 = 13.8986287281 kΩ T = RC = 13898.6287281 * 3.5E-15 = 4.86452005484e-11 t = 4.86452005484e-11 * 1.94591014906 = 9.46591894502e-11 = 94.6591894502 ps

OpenStudy (anonymous):

that is the same calculation as above, just different values. Still, I don't know whats wrong with it.

OpenStudy (anonymous):

I did the same reasoning but with different values ​​and also not getting the correct answer